can you help me evaluate the expression or let me know if this is correct
a=6 and b=1
5(a+4b) + 3(2a-7b)
30+20+36-21=65
is that correct?

Answer by Alan3354(69443)

a=6 and b=1
5(a+4b) + 3(2a-7b)
= 5(6+4) + 3(12-7)
= 5*10 + 3*5
= 65
------
30+20+36-21=65